OpenStudy (anonymous):

Given the parent functions f(x) = 3x + 2 and g(x) = 5x − 10, what is g(x) − f(x)?

OpenStudy (anonymous):

g(x) is the same thing as 5x-10 f(x) is the same thing as 3x+2 The above means that if you're looking for g(x)-f(x) you simply replace g(x) and f(x) with what they are.

OpenStudy (anonymous):

i tried using the formula f(x) - g(x) ... but when i solved it did not fit any of my choices which are g(x) − f(x) = 5x − 8 − 3x g(x) − f(x) = 5x − 12 − 3x g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x + 12 − 5x

OpenStudy (anonymous):

by solving i got -9, but that's not what they want

OpenStudy (anonymous):

You don't have to solve it

OpenStudy (anonymous):

Ok, I'll try to simplify what's being said even more. f(x)=3x+2 g(x)=5x-10 g(x)-f(x)=?

OpenStudy (anonymous):

5x - 10 - 3x + 2

OpenStudy (anonymous):

Don't forget that -f(x) means -(3x+2)

OpenStudy (anonymous):

g(x)-f(x)=(5x-10)-(3x+2)=5x-10-3x-2=2x-12

OpenStudy (anonymous):

SHIRIN74: Read the Code of Conduct please.

OpenStudy (anonymous):

but look at my answer choices ... i'm still not understanding

OpenStudy (anonymous):

Multiple choice questions are stupid so you'll have to rearrange your answer to fit the answers.

OpenStudy (anonymous):

Noliec :what?i'm completely sure about mu answer

OpenStudy (anonymous):

Yes, you're never supposed to just give the answer.

OpenStudy (anonymous):

hirin can you break it down for me

OpenStudy (anonymous):

What you wrote was almost correct mgarrick77

OpenStudy (anonymous):

somebody is still incorrect ` my answer choices are g(x) − f(x) = 5x − 8 − 3x g(x) − f(x) = 5x − 12 − 3x g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x + 12 − 5x

OpenStudy (anonymous):

You have to remember though, like I said before that -f(x) is not -3x+2

OpenStudy (anonymous):

-(3x+2) is the same thing as -3x-2

OpenStudy (anonymous):

the 2nd choice is correct. g(x)-f(x)= 5x-10-(3x+2) =5x-10-3x-2 =5x-10-2-3x =5x-12-3x Hope this helps!

OpenStudy (anonymous):

lol that's what i was asking ... was i suppose to add 10+2

OpenStudy (anonymous):

okay the answer is g(x) − f(x) = 5x − 12 − 3x

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

5x-3x=2x did u know that?

OpenStudy (anonymous):

yes shirin i knew that

OpenStudy (anonymous):

can you guys continue to help me ? i want to make sure i got it

OpenStudy (anonymous):

okay it's done

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Given the function f(x) = 0.3(4)x, what is the value of f−1(6)? 0.360 1.242 2.161 2.312 i guessing i suppose to plug in ... but when i took the numbers it didn't quite work

OpenStudy (anonymous):

i have a Ti-84 calculator ... i tried to plug it in

OpenStudy (anonymous):

You have that: f(x)=3x+2 g(x)=5x-10 You're looking for g(x)-f(x), which is the same thing as (5x-10)-(3x+2), which is the same thing as 5x-10-3x-2, try to see how you can reformulate this to your question.

OpenStudy (anonymous):

can we say: f(x) = 0.3(4)x=1.2x?

OpenStudy (anonymous):

That's a better start than initially Shirin74! ;)

OpenStudy (anonymous):

noliec:the answer of that question was found and now it's finished we are thrying to answer the new one

OpenStudy (anonymous):

(in terms of following the code of conduct and actually emphasizing on learning people things) =)

OpenStudy (anonymous):

The new question is one of the worst questions I've seen in academic history. The use of a calculator in math should be prohibited and problems designed so that you don't need one.

OpenStudy (anonymous):

my new question is up their

OpenStudy (anonymous):

so what should we do?

OpenStudy (anonymous):

i tried by pluging in the f^-1(6)